OpenStudy (anonymous):

@Mertsj Graph the Ellipse and locate the foci x^2/25+y^2/64=1

OpenStudy (anonymous):

We got the graph we just need the foci

OpenStudy (mertsj):

a^2-b^2=c^2

OpenStudy (mertsj):

Find c because the foci are c units to the right and c units to the left of the center.

OpenStudy (anonymous):

sqrt 39?

OpenStudy (mertsj):

yes. And I just noticed that the major axis is vertical so the foci are c units above and c units below the center.

OpenStudy (mertsj):

So the foci are: \[(0,\pm \sqrt{39})\]

OpenStudy (anonymous):

he got that

OpenStudy (anonymous):

Try drawing a picture if you can. Try that first and see if you can "see" the solution and then move on the the algebra.

OpenStudy (mertsj):

He was right.

OpenStudy (mertsj):

Tell him "good job"

OpenStudy (anonymous):

question is comleted